OpenStudy (anonymous):

3. How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

\[3x + 7y = –10\\ 15x + 35y = –45\] multiply the first equation by \(5\) and you get \[15x + 35y = –50 \\ 15x + 35y = –45\] both the left hand sides are identical, but the right hand sides are different, so you cannot solve it. i.e. the lines are parallel